Grade 11Mechanics

a grass hopper can jump maximum distance 1.6m . it spends negligible time on ground . how far can it go in 10×2^1/2

prarthna
8 Years agoGrade 11

1 Answer

Venkat
8 Years ago
Note thta i have solved for two different values of time as the qn is not clear about time.
For 10s and 5 secs, the solutions is presented below.
 
Horizontal distance jumped will be maximu for 45 degrees, Hence,
\\R_{max}=\frac{u^2}{g} \\1.6=\frac{u^2}{10} \\u=4m/s
 
Horziontal component is,
ucos45=4\times 1/\sqrt2=2\sqrt{2}
 
Max distance covered in 10 secs is ,
u(cos45)t=2\sqrt{2}\times 10=20\sqrt{2}
 
Max distance covered in 5 secs is ,
u(cos45)t=2\sqrt{2}\times 5=10\sqrt{2}